OpenStudy (anonymous):

someone HELP 6x(32-90)9-8=1 what is the value of x

OpenStudy (anonymous):

try multiplying random numbers and do the rest of the math untill it matches

OpenStudy (anonymous):

thank u :)

pooja195 (pooja195):

6x(32−90)(9)−8=1 −3132x+−8=1 −3132x−8=1 −3132x−8+8=1+8 −3132x=9 Divide both sides by -3132

OpenStudy (anonymous):

wow thanks that was hard @pooja195

pooja195 (pooja195):

np :) your answer should be a fraction

OpenStudy (anonymous):

thank you :)